### v4

# A representation for the Fremlin projective tensor product between Banach lattices: a topological point of view

Abstract.
For an arbitrary topological space $X$, assume that $S(X)$ is the vector lattice of all equivalence classes of real-valued continuous functions on open dense subsets of $X$; it is a laterally complete vector lattice but not a normed lattice, certainly. Nevertheless, we can have the extended unbounded norm topology ($un$-topology) on it. On the other hand, by a remarkable result of Wickstead, there exists a representation approach for every Archimedean vector lattice $E$ in terms of $S(X)$-spaces. In this paper, we show that this representation is order continuous and when $E$ is order complete, it coincides with the known Maeda-Ogasawara representation. Moreover, when $E$ is a Banach lattice, by consideration of the $un$-topology on $E$ and the extended $un$-topology on $S(X)$, we show that this representation is, in fact, a homeomorphism. With the aid of this topological attitude, we establish a representation theorem (in fact a homeomorphism) for the Fremlin projective tensor product between Banach lattices, in terms of $S(X)$-spaces, as well.
Key words and phrases:
Unbounded order convergence, order continuous Banach lattice, unbounded norm convergence, Fremlin projective tensor product, Fremlin tensor product.
2020 Mathematics Subject Classification:
Primary: 46M05. Secondary: 46A40.
August 4, 2025
1. Motivation and introduction
Let us start with some motivation. Assume that $E$ is an Archimedean vector lattice. There are two important and significant representations of $E$ in terms of some functions spaces. The first one is the known Maeda-Ogasawara representation theorem which states that $E$ can be considered as an order dense vector sublattice of some $C^{\infty}(\Omega)$-space, where $\Omega$ is an extremally disconnected compact Hausdorff space; in fact, $C^{\infty}(\Omega)=E^{u}$, the universal completion of $E$ (see [1, Chapter 7] for a comprehensive explanation). This approach is practical and useful; however, there is one problem, here. $C^{\infty}(X\times Y)$ may not be a vector space even if $X$ and $Y$ are extremally disconnected compact Hausdorff spaces (see [5] for more details). This problem causes difficulty while we are dealing with tensor products. Let us explain more. By the remarkable Kakutani theroem every Archimedean vector lattice $X$ with an order unit is a norm and order dense sublattice of a $C(K)$-space for some compact Hausdorff space $K$. Furthermore, when $X$ and $Y$ are Archimedean vector lattices with order units and with the representations $C(K_{1})$ and $C(K_{2})$ ($K_{1}$ and $K_{2}$ compact Hausdorff spaces), respectively, the Fremlin tensor product $X\overline{\otimes}Y$ is norm and order dense in $C(K_{1}\times K_{2})$; which is certainly a vector lattice. Now, if we want to develop such theory for arbitrary vector lattices (with $C^{\infty}(\Omega)$-spaces instead of $C(K)$-spaces), we must overcome the mentioned difficulty.
Buskes and Wickstead in [5] solved this problem by considering the space $S(X)$ for any topological space $X$. In fact, $S(X)$ consists of all equivalence classes of continuous real-valued functions defined on open dense subsets of $X$; it is an Archimedean vector lattice under pointwise lattice and algebraic operations. In this case, we can have a representation for the Fremlin tensor product of general Archimedean vector lattices in terms of $S(X)$-spaces (for a complete context on this approach, see [5, 17]). Recently, Wickstead in [15], established a representation theorem for Archimedean vector lattices in terms of $S(X)$-spaces.
In this paper, we show that this representation is order continuous. Furthermore, when $E$ is order complete, we prove that $S(X)$ is order complete and also $X$ is extremally disconnected. Therefore, $S(X)$ and $C^{\infty}(X)$ agree. On the other hand, since $S(X)$ is always laterally complete, it can not be a normed lattice; however, we can have unbounded norm topology ($un$-topology) on it; by the extended $un$-topology considered in [13]. These points motivate us to investigate the representation posed by Wickstead for vector lattices, for the case when $E$ is also a Banach lattice. In fact, we show that when $E$ is an order continuous Banach lattice, there is a representation of $E$ into a $S(X)$-space that is also a homeomorphism (we equip $E$ with the $un$-topology while $S(X)$ enjoys the extended $un$-topology). Furthermore, with using this attitude, we are able to establish a homeomorphism representation for the Fremlin projective tensor product between Banach lattices in terms of $S(X)$-spaces, as well. In the sequel, we recall some notes regarding unbounded convergences as well as the Fremlin tensor products between vector and Banach lattices.
2. preliminaries
2.1. unbounded convergences
[…]
2.2. Fremlin tensor product
In this part, we recall some notes about the Fremlin tensor product between vector lattices and Banach lattices. For more details, see [7, 8]. Furthermore, for a comprehensive, new and interesting reference, see [16]. Moreover, for a short and nicely written exposition on different types of tensor products between Archimedean vector lattices, see [10].
Assume that $E$ and $F$ are Archimedean vector lattices. In 1972, Fremlin constructed a tensor product $E\overline{\otimes}F$ that is an Archimedean vector lattice such that the algebraic tensor product $E\otimes F$ is a vector subspace of $E\overline{\otimes}F$ so that it is an ordered vector subspace in its own right. Moreover, the vector sublattice in $E\overline{\otimes}F$ generated by $E\otimes F$ is the whole of $E\overline{\otimes}F$. Therefore, we conclude that every element of $E\overline{\otimes}F$ can be considered as a finite supremum and finite infimum of some elements of $E\otimes F$.
Now, let $E$ and $F$ be Banach lattices. Fremlin in [8] constructed a tensor product $E\widehat{\otimes}F$ that is a Banach lattice. In fact, $E\widehat{\otimes}F$ is the norm completion of $E\otimes F$ with respect to the projective norm: for each $u=\Sigma_{i=1}^{n}x_{i}\otimes y_{i}\in E\otimes F$:
$\|u\|_{|\pi|}=\sup\{|\Sigma_{i=1}^{n}\phi(x_{i},y_{i})|:\phi\hskip 7.11317pt\textit{is a bilinear form on}\hskip 9.95863ptE\times F\textit{and}\hskip 7.11317pt\|\phi\|\leq 1\}.$
Furthermore, $E\overline{\otimes}F$ can be considered as an norm-dense vector sublattice of $E\widehat{\otimes}F$. Moreover, the projective norm, .$\|.\|_{|\pi|}$, on $E\widehat{\otimes}F$ is a cross norm; that is for every $x\in E$ and for every $y\in F$, we have $\|x\otimes y\|_{|\pi|}=\|x\|\|y\|$. For a comprehensive explanation and also different properties of related to these tensor products, see [7, 8].
3. main results
First, we start with Archimedean vector lattices and $S(X)$-spaces. It is shown in [15, Proposition 2.5] that $S(X)$-spaces are laterally complete. Now, we proceed to see whether or not this space is order complete.
Lemma 1.
Suppose $X$ is a completely regular topological space. Then $S(X)$ is order complete if and only if $X$ is extremally disconnected.
Proof.
Suppose $X$ is extremally disconnected. By [17, Lemma 1], $S(X)$ is lattice isomorphic to $C^{\infty}(X)$ so that $S(X)$ is order complete. For the converse, suppose $S(X)$ is order complete. Using [15, Proposition 2.5], convinces us that it is universally complete. Therefore, there exists an extremally disconnected compact Hausdorff $Q$ such that $S(X)$ is lattice isomorphic to $C^{\infty}(Q)$; assume that $\phi$ is the desired lattice isomorphism between $S(X)$ and $C^{\infty}(Q)$. On the other hand, by [17, Lemma 1] again, we can identify $C^{\infty}(Q)$ with $S(Q)$. Note that both $S(X)$ and $S(Q)$ have weak units $\textbf{1}_{X}$ and $\textbf{1}_{Q}$, respectively. Thus, by [15, Theorem 4.1 and Theorem 4.2], there is a continuous function $\pi:Q\to X$ such that $\phi(f)=fo\pi$. This shows that the restriction $\phi$ to $C(X)$ maps $C(X)$ into $C(Q)$ and is still a lattice isomorphism. Thus $X$ and $Q$ are homeomorphic so that $X$ is extremally disconnected, as well. ∎
Suppose $E$ is an Archimedean vector lattice with a weak unit $u$. By [15, Theorem 3.3], there exists a compact Hausdorff space $X$ and a lattice isomorphism $T:E\to S(X)$ such that $T(u)={\textbf{1}_{X}}$ and the norm closure (with the sup-norm) of the image of the ideal generated by $u$ in $E$, $E_{u}$, is $C(X)$. In the following, we show that $T$ preserves $uo$-convergence.
Lemma 2.
Suppose $E$ is an Archimedean vector lattice $E$ with a weak unit $u$ and $T$ is the representation of $E$ as described in [15, Theorem 3.3]. Then $T$ is both $uo$-continuous and order continuous. Moreover, $T^{-1}$ is also $uo$-continuous.
Proof.
By [4, Theorem 5.2], $uo$-continuity and order continuity of $T$ are equivalent. Moreover, suppose $E_{u}$ is the ideal generated by $u$ in $E$ and $I_{\textbf{1}_{X}}$ is the ideal in $S(X)$ generated by the constant function $\textbf{1}_{X}$. It is easy to see that $T(E_{u})$ is a vector sublattice of $I_{\textbf{1}_{X}}$. Note that $T(E_{u})$ is norm dense in $C(X)$ so that it is order dense by [7, Lemma 1.2]. On the other hand, by [17, Lemma 5], $C(X)$ is order dense in $S(X)$. Therefore, we see that $T(E_{u})$ is order dense in $S(X)$. This implies that $T(E)$ is order dense in $S(X)$, as well. By considering [9, Corollary 3.5] and [9, Theorem 3.2], we conclude that $uo$-continuity in $E$ and $S(X)$ reduces to order continuity of the restriction of $T$ to $E_{u}$ onto $T(E_{u})$. Now, by [2, Theorem 2.21], we have the desired result. Note that by [9, Theorem 3.2], $T(E)$ is a regular Riesz subspace of $S(X)$ so that by [9, Theorem 3.2], $uo$-convergence has the same meaning in both $T(E)$ and $S(X)$. Therefore, $T^{-1}$ is also $uo$-continuous. ∎
Now, we extend this result to the case when the Archimedean vector lattice does not have weak units. Before that, we have a simple observation.
Remark 3.
Suppose $(X_{\alpha})$ is a family of topological spaces and $\sqcup_{\alpha}X_{\alpha}$ is the disjoint unions of them. Then, we can identify $S(\sqcup_{\alpha}X_{\alpha})$ and the Cartesian product $\prod_{\alpha}S(X_{\alpha})$, topologically and ordering. More precisely, for each family $(f_{\alpha})\subseteq\prod_{\alpha}S(X_{\alpha})$, there exists a necessarily unique $f\in S(\sqcup_{\alpha}X_{\alpha})$ such that $f=(f_{\alpha})_{\alpha}$. Furthermore, since the ordering on the Cartesian product is assumed to be componentwise, it is easily seen that $|f|=(|f_{\alpha}|)_{\alpha}$. Moreover, it can be verified that if $Y_{\alpha}$ is an order dense vector sublattice of $S(X_{\alpha})$, then $\prod_{\alpha}Y_{\alpha}$ is an order dense vector sublattice of $\prod_{\alpha}S(X_{\alpha})=S(\sqcup_{\alpha}X_{\alpha})$.
Theorem 4.
Suppose $E$ is an Archimedean vector lattice and $T$ is the representation of $E$ as described in [15, Corollary 3.6]. Then, $T$ is both $uo$-continuous and order continuous. Moreover, $T^{-1}$ is also $uo$-continuous.
Proof.
Again by [4, Theorem 5.2], $uo$-continuity and order continuity of $T$ are equivalent. We show that $T$ is order continuous. We have a decomposition of $E$ into direct sums of pairwise disjoint bands $(B_{\alpha})_{\alpha\in I}$, each of them has a weak unit, namely, $x_{\alpha}$ (by [11, Theorem 28.5]). By Lemma 2, there are compact Hausdorff spaces $X_{\alpha}$ and order continuous lattice isomorphisms $T_{\alpha}:B_{\alpha}\to S(X_{\alpha})$. Put $X=\sqcup_{\alpha}X_{\alpha}$, the Hausdorff space of the disjoint unions of $X_{\alpha}^{\prime}s$. By using [15, Corollary 3.6], we see that $T(x)=(T_{\alpha}(y_{\alpha}))_{\alpha}$, in which $x=\vee_{\alpha}y_{\alpha}$ and $y_{\alpha}\in B_{\alpha}$. By [9, Theorem 3.2], it is enough to show that $T(E)$ is a regular vector sublattice of $S(\sqcup_{\alpha}X_{\alpha})$. Note that for each $\alpha$, by Lemma 2, $T_{\alpha}(B_{\alpha})$ is order dense in $S(X_{\alpha})$. Therefore, form Remark 3 the conclusion follows.
Among the proof of Theorem 4, we see that $E$ (identifying with $T(E)$) is an order dense vector sublattice of $S(X)$. So, by using [2, Theorem 2.31], the following result follows.
Corollary 5.
Suppose $E$ is an order complete vector lattice and $T:E\to S(X)$ is its representation as described in Theorem 4. Then, $E$ can be considered as an ideal of $S(X)$.
Now, as an application, we present a representation approach for the Fremlin tensor product between Archimedean vector lattices. For details see [7]; for a new and more illustrative approach, see [16].
Corollary 6.
Suppose $E$ and $F$ are Archimedean vector lattices. Moreover, assume that $S(X)$ and $S(Y)$ are the corresponding representations of $E$ and $F$, respectively as described in Theorem 4. Then, there exists a similar representation for the Fremlin tensor product $E\overline{\otimes}F$ in $S(X\times Y)$, as well.
Proof.
By Theorem 4, there are Hausdorff spaces $X$ and $Y$ and order continuous lattice isomorphisms $T:E\to S(X)$ and $S:F\to S(Y)$. Consider the bi-injective lattice bimorphism defined via $(x,y)\to T(x)\otimes S(y)$ from $E\times F$ into $S(X\times Y)$; it induces a lattice homomorphism $T\otimes S:E\overline{\otimes}F\to S(X\times Y)$ defined via $(T\otimes S)(x\otimes y)=T(x)\otimes S(y)$. We show that $T\otimes S$ is also one-to-one. First, assume that $T(x\otimes y)=T(x)\otimes S(y)=0$. By [5, Proposition 3.1], we see that $T(x)=0$ or $S(y)=0$. Therefore, $x=0$ or $y=0$ so that $x\otimes y=0$. Assume that $0\neq u\in(E\overline{\otimes}F)_{+}$ with $(T\otimes S)(u)=0$, by [7, Theorem 4.2 (4)], there exist $x_{0}\in E_{+}$ and $y_{0}\in F_{+}$ with $0<x_{0}\otimes y_{0}\leq u$ so that $T(x_{0})\otimes S(y_{0})=0$. Therefore, by the previous part, $x_{0}\otimes y_{0}=0$ which is a contradiction. Thus, $T(E)$ and $S(F)$ can be considered as order dense vector sublattices of $S(X)$ and $S(Y)$, respectively. By [17, Lemma 7 and Lemma 8], we conclude that $T(E)\overline{\otimes}S(F)$ can be considered as an order dense vector sublattice of $S(X\times Y)$. It can be verified that $(T\otimes S)(E\overline{\otimes}F)$ is an order dense vector sublattice of $T(E)\overline{\otimes}S(F)$. Now, [9, Theorem 3.2] convinces that $uo$-continuity can be transferred between $E\overline{\otimes}F$ and $S(X\times Y)$ via $T\otimes S$. ∎
It is known that the universal completion of an Archimedean vector lattice $E$, $E^{u}$, can be identified with a $C^{\infty}(\Omega)$-space, in which, $\Omega$ is an extremally disconnected compact Hausdorff topological space. On the other hand, we can have the lateral completion of $E$, $E^{\lambda}$, in a similar manner: the intersection of all laterally complete Archimedean vector lattices that contain $E$ as a vector sublattice. By [1, Page 213, Exercise 10], $(E^{\delta})^{\lambda}=(E^{\lambda})^{\delta}=E^{u}$, in which, $E^{\delta}$ is the order completion of $E$. It is interesting to know that the lateral completion also has the form of a $S(X)$-space.
Corollary 7.
Suppose $E$ is an order complete vector lattice and $T$ is the lattice isomorphism from $E$ into $S(X)$ as described in [15, Corollary 3.6]. Then, $E^{\lambda}=E^{u}=S(X)$ and $X$ is extremally disconnected.
Proof.
By Theorem 4, $T$ is an order continuous lattice isomorphism. Since by [15, Proposition 2.5], $S(X)$ is laterally complete and by the assumption, $E$ is order complete, by [2, Theorem 2.32], $T$ can have an extension to an order continuous lattice isomorphism (denoted by $T$, again) from $E^{\lambda}=E^{u}$ into $S(X)$. By using Corollary 5 and also [9, Theorem 3.2], we can see that $E^{u}$ is an order dense vector sublattice of $S(X)$. By [1, Theorem 7.15], $E^{\lambda}=E^{u}$ and $S(X)$ are in fact the same. Therefore, by Lemma 1, $X$ is extremally disconnected. ∎
Now, we are going to establish the main results of the note. First, we need some preliminaries.
Suppose $X$ is a topological space. Note that $S(X)$ can not be a normed lattice; nevertheless, by the extended $un$-convergence procedure described in [13], we can have $un$-topology on $S(X)$: suppose $Y$ is an Archimedean vector lattice and $E$ is a normed lattice which is an ideal in $Y$, as well. For a net $(y_{\alpha})\subseteq Y$, we say $y_{\alpha}$ is unbounded norm convergent ($un$-convergent) to $y\in Y$ if for every $x\in E_{+}$, $\||y_{\alpha}-y|\wedge x\|\rightarrow 0$. This convergence induces a topology on $Y$ which is called ”the extended $un$-topology” on $Y$ induced by $E$. However, this convergence is dependent on the ideal $E$; moreover, it may not be Hausdorff, in general. The good news is that the extended $un$-topology induced by an ideal is Hausdorff if and only if the ideal is order dense (see [13, Proposition 1.4]).
[…]
Theorem 8.
Suppose $E$ is an order continuous Banach lattice. Then, there exists a Hausdorff topological space $X$ and a lattice isomorphism homeomorphism $T:(E,un)\to(S(X),\tau_{E})$, in which $S(X)$ is equipped with the extended $un$-topology induced by $E$.
Proof.
First, assume that $E$ has a weak unit (quasi-interior point) $u$ and note that we can identify the image of $E$ under $T$ ($T(E)$) with $E$ (topologically and ordering). Consider order continuous lattice isomorphism $T:E\to S(X)$ as described in Theorem 4. $T(E)$ can be considered an an order dense vector sublattice of $S(X)$ so that by [2, Theorem 2.31], it is an ideal. So, we can consider the extended $un$-topology on $S(X)$ induced by $E$. Since $E$ is order continuous, the induced topology on $S(X)$ is linear; that is $S(X)$ is a locally solid vector lattice. By [1, Theorem 5.19], $T$ is continuous. Furthermore, the extended $un$-topology is also metrizable by [13, Theorem 3.3]. So, by [2, Theorem 7.55], it is also Lebesgue.
Now, suppose $(f_{\alpha})\subseteq E$ is $un$-null so that $\|f_{\alpha}\wedge u\|\rightarrow 0$. Therefore $\|T(f_{\alpha})\wedge\textbf{1}_{X}\|\rightarrow 0$. Now, [13, Proposition 3.1] implies that $T(f_{\alpha})\xrightarrow{un-E}0$ in $S(X)$. Now assume that a net $(g_{\alpha})\subseteq T(E)$ is $un$-null. There exists a net $(f_{\alpha})\subseteq E$ with $T(f_{\alpha})=g_{\alpha}$. Since $g_{\alpha}\xrightarrow{un}0$, we conclude that $\|g_{\alpha}\wedge\textbf{1}_{X}\|_{T(E)}=\|T(f_{\alpha}\wedge u)\|_{T(E)}=\|f_{\alpha}\wedge u\|_{E}\rightarrow 0$. Thus, $f_{\alpha}\wedge u\xrightarrow{un}0$ so that $(f_{\alpha})$ is $un$-null since $u$ is a quasi-interior point.
Now, we proceed with the general case. Suppose $E$ is an order continuous Banach lattice. By [14, Proposition 1.a.9], it possesses a dense band decomposition. More precisely, there exists a pairwise disjoint family of bands B such that every band $B_{\alpha}$ in B has a weak unit (quasi-interior point, namely, $x_{\alpha}$) and $E$ is the closure of the direct sums of all elements in B. By the former case, there exist compact Hausdorff spaces $X_{\alpha}$ and lattice isomorphisms $T_{\alpha}:B_{\alpha}\to S(X_{\alpha})$ that are $un$-homeomorphisms. By using [2, Thoerem 2.14], we conclude that for each $_{\alpha_{1}},B_{\alpha_{2}}\in\textbf{B}$ B, $T_{\alpha_{1}}(B_{\alpha_{1}})\wedge T_{\alpha_{2}}(B_{\alpha_{2}})=0$. Now, define the lattice isomorphism $T:E\to S(\sqcup_{\alpha}X_{\alpha})$ defined via $T(x)=T(\Sigma_{\alpha}y_{\alpha})=T(\bigvee_{\alpha}y_{\alpha})=\bigvee_{\alpha}(T_{\alpha}(y_{\alpha}))$. Note that $S(\sqcup_{\alpha}X_{\alpha})$ can be identified with $\prod_{\alpha}S(X_{\alpha})$ by Remark 3.
For each $\alpha$, assume that $P_{\alpha}$ is the natural band projection from $E$ onto $B_{\alpha}$; $B_{\alpha}$ is an order continuous Banach lattice with a quasi-interior point that induces a Hausdorff locally solid Lebesgue topology $\tau_{\alpha}$ on $S(X_{\alpha})$ by the former case. Now, consider the product topology $\tau$ on $S(\sqcup_{\alpha}X_{\alpha})=\prod_{\alpha}S(X_{\alpha})$; it is a Hausdorff locally solid topology by [1, Theorem 2.20] and also Lebesgue by [1, Theorem 3.11]. On the other hand, we have the extended $un$-topology $\tau_{E}$ on $S(\sqcup_{\alpha}X_{\alpha})$ that is Hausdorff (since $E$ is order dense in $S(X)$), locally solid (since $E$ is order continuous). On the other hand, by Proposition 7 and also by using [13, Theorem 6.7], $S(\sqcup_{\alpha}X_{\alpha})$ is $un$-complete (with respect to the induced topology $\tau_{E}$). Furthermore, by [13, Proposition 9.1], $(S(\sqcup_{\alpha}X_{\alpha}),\tau_{E})$ satisfies the pre-Lebesgue property so that by [1, Theorem 3.26], $\tau_{E}$ is also Lebesgue. Therefore, by [1, Theorem 7.53], we have $\tau=\tau_{E}$.
[…]
For the converse, assume that a net $T(x_{\beta})\subseteq T(E)$ is $\tau$-null. Assume that $x_{\beta}=(y^{\beta}_{\alpha})_{\alpha}$, in which $y^{\beta}_{\alpha}\in B_{\alpha}$. Therefore, $T_{\alpha}(y^{\beta}_{\alpha})=T_{\alpha}P_{\alpha}(x_{\beta})\xrightarrow{un}0$ in $S(X_{\alpha})$. By the former case, $P_{\alpha}(x_{\beta})=y^{\beta}_{\alpha}\xrightarrow{un}0$ in $B_{\alpha}$. Now, [12, Theorem 4.12], convinces us that $x_{\beta}\xrightarrow{un}0$ in $E$ as claimed.
As an application, we establish a $un$-homeomorphism representation for the Fremlin projective tensor product of Banach lattices; for more details, see [8]. Before that, we show that quasi-interior points can be preserved by the Fremlin projective tensor product $E\widehat{\otimes}F$.
Proposition 9.
Suppose $E$ and $F$ are Banach lattices with quasi-interior points. Then, the Fremlin projective tensor product $E\widehat{\otimes}F$ has a quasi-interior point, as well.
Proof.
Assume that $E$ has a quasi-interior $x_{0}$ and $F$ possesses a quasi-interior point $y_{0}$. We show that $x_{0}\otimes y_{0}$ is a quasi-interior point for $E\widehat{\otimes}F$. By [2, Theorem 4.85], it is enough to show that for every $0<f\in(E\widehat{\otimes}F)^{\prime}$, $f(x_{0}\otimes y_{0})>0$. By [8, 1(A) d], $(E\widehat{\otimes}F)^{\prime}=B^{r}(E\times F)$ where $B^{r}(E\times F)$ is the Banach lattice of all bounded regular bilinear forms on $E\times F$. There exist $0\neq x_{1}\in E_{+}$ and $0\neq y_{1}\in F$ with $f(x_{1},y_{1})>0$ so that $f(x_{1},y_{0})\neq 0$ since the restriction $f$ to $x_{1}$, is a non-zero positive functional on $Y$. On a contrary, assume that $f(x_{0},y_{0})=0$. Note that $x_{1}\wedge nx_{0}\rightarrow x_{1}$ so that $f(x_{1}\wedge nx_{0},y_{0})\rightarrow f(x_{1},y_{0})$, since $f$ is continuous. Note that $f(x_{1}\wedge nx_{0},y_{0})\leq nf(x_{0},y_{0})=0$ which is a contradiction. ∎
Suppose $E$ is a Banach lattice. Recall that $E$ possesses a dense band decomposition if there exists a family B of pairwise disjoint projection bands in $E$ such that the linear span of all of the bands in B is norm dense in $E$. For more details see [12, Section 4.1]. For example by [14, Proposition 1.a.9], every order continuous Banach lattice possesses a dense band decomposition; see also [12, Theorem 4.11]. In the following, we show that if Banach lattices $E$ and $F$ have dense band decompositions, then, so is the Fremlin projective tensor product $E\widehat{\otimes}F$.
Lemma 10.
Suppose $E$ and $F$ are Banach lattices such that the Fremlin tensor product $E\overline{\otimes}F$ is order complete. Moreover, assume that $\textbf{B}=(B_{\alpha})_{\alpha\in I}$ and $\textbf{C}=(C_{\beta})_{\beta\in J}$ are dense band decompositions $E$ and $F$, respectively. Then, the collection $\textbf{A}=\{B_{\alpha}\widehat{\otimes}C_{\beta};B_{\alpha}\in\textbf{B},C_{\beta}\in\textbf{C}\}$ forms a dense band decomposition for $E\widehat{\otimes}F$.
Proof.
First, observe that by [10, Proposition 3.9], both $E$ and $F$ are order complete. Put ${\textbf{A}_{0}}=\{B_{\alpha}\overline{\otimes}C_{\beta};B_{\alpha}\in\textbf{B},C_{\beta}\in\textbf{C}\}$. By [3, Theorem 5.8], we see that each element of ${\textbf{A}_{0}}$ is a projection band in $E\overline{\otimes}F$. By [1, Theorem 2.48], the elements of A are also projection bands in $E\widehat{\otimes}F$. Moreover, the elements of A are pairwise disjoint. We claim that A is a dense band decomposition for $E\widehat{\otimes}F$. We use [12, Lemma 4.10]. Note that the elements of A are pairwise disjoint. For each $\alpha,\alpha^{\prime},\beta,\beta^{\prime}$, we have $(B_{\alpha}\otimes C_{\beta})\wedge(B_{\alpha^{\prime}}\otimes C_{\beta^{\prime}})=0$. Otherwise, for each non-zero positive $u\in B_{\alpha}\otimes C_{\beta})\wedge(B_{\alpha^{\prime}}\otimes C_{\beta^{\prime}})$, by [8, 1(A) d], we can find $x_{0}\in{B_{\alpha}}_{+}$, $x_{1}\in{B_{\alpha^{\prime}}}_{+}$, $y_{0}\in{C_{\beta}}_{+}$ and $y_{1}\in{C_{\beta^{\prime}}}_{+}$ with $u\leq x_{0}\otimes y_{0}$ and $u\leq x_{1}\otimes y_{1}$. So,
$u\leq(x_{0}\otimes y_{0})\wedge(x_{1}\otimes y_{1})\leq(x_{0}\wedge x_{1})\otimes(y_{0}\vee y_{1})=0,$
that is a contradiction. Now, it is routine to check that if for two subsets $A,B$ in a normed lattice $A\wedge B=0$, then, $\overline{A}\wedge\overline{B}=0$. So, the elements of A are pairwise disjoint. We need to characterize band projections for elements of A. Fix $\alpha\in I$ and $\beta\in J$. Assume that $S_{\alpha,\beta}$ is the corresponding band projection from $E\widehat{\otimes}F$ onto $B_{\alpha}\widehat{\otimes}C_{\beta}$. Also, assume that $P_{\alpha}$ and $Q_{\beta}$ are corresponding band projections onto $B_{\alpha}$ and $C_{\beta}$, respectively. By considering the lattice bimorphism $\sigma:E\times F\to B_{\alpha}\overline{\otimes}C_{\beta}$ defined by $\sigma(x,y)=P_{\alpha}(x)\otimes Q_{\beta}(y)$ and using [8, 1A(b)], there exists a lattice homomorphism $P_{\alpha}\otimes Q_{\beta}:E\overline{\otimes}F\to B_{\alpha}\overline{\otimes}C_{\beta}$ via $(P_{\alpha}\otimes Q_{\beta})(x\otimes y)=P_{\alpha}(x)\otimes Q_{\beta}(y)$. We claim that $P_{\alpha}\otimes Q_{\beta}=S_{\alpha,\beta}$ on $E\overline{\otimes}F$ and so on $E\widehat{\otimes}F$ by taking a norm completion.
Assume that $x\in E$ and $y\in F$. We can write $x\otimes y=u_{\alpha,\beta}+v_{\alpha,\beta}$ in which, $u_{\alpha,\beta}\in B_{\alpha}\overline{\otimes}C_{\beta}$ and $v_{\alpha,\beta}\in{(B_{\alpha}\overline{\otimes}C_{\beta})}^{d}$ and this representation is unique. On the other hand, we can also write $x=r_{\alpha}+{r_{\alpha}}^{d}$ and $y=w_{\beta}+{w_{\beta}}^{d}$, in which $r_{\alpha}\in B_{\alpha}$, ${r_{\alpha}}^{d}\in{B_{\alpha}}^{d}$, $w_{\beta}\in C_{\beta}$ and ${w_{\beta}}^{d}\in{C_{\beta}}^{d}$. Therefore, we have
$x\otimes y=(r_{\alpha}+{r_{\alpha}}^{d})\otimes(w_{\beta}+{w_{\beta}}^{d})=r_{\alpha}\otimes w_{\beta}+{r_{\alpha}}^{d}\otimes w_{\beta}+r_{\alpha}\otimes{w_{\beta}}^{d}+{r_{\alpha}}^{d}\otimes{w_{\beta}}^{d}.$
Note that $r_{\alpha}\otimes w_{\beta}\in B_{\alpha}\overline{\otimes}C_{\beta}$ and ${r_{\alpha}}^{d}\otimes w_{\beta}+r_{\alpha}\otimes{w_{\beta}}^{d}+{r_{\alpha}}^{d}\otimes{w_{\beta}}^{d}\in(B_{\alpha}\overline{\otimes}C_{\beta})^{d}$. Thus, $u_{\alpha,\beta}=r_{\alpha}\otimes w_{\beta}$ by uniqueness of the representation. So, $S_{\alpha,\beta}(x\otimes y)=u_{\alpha,\beta}=r_{\alpha}\otimes w_{\beta}=P_{\alpha}(x)\otimes Q_{\beta}(y)$. Therefore, $P_{\alpha}\otimes Q_{\beta}=S_{\alpha,\beta}$ on $E\otimes F$. Every element of $E\overline{\otimes}F$ is a finite suprema and a finite infima of some elements of $E\otimes F$. Since $S_{\alpha,\beta}$ is a band projection, it is order continuous lattice homomorphism so that $P_{\alpha}\otimes Q_{\beta}=S_{\alpha,\beta}$ on $E\overline{\otimes}F$ and so on $E\widehat{\otimes}F$ by an extension that is also a band projection, as well.
Now, suppose $x\in E$ and $y\in F$ and also $\varepsilon>0$ is arbitrary. By [12, Lemma 4.10], we can find indices $\{\alpha_{1},\ldots,\alpha_{r}\}$ and also $\{\beta_{1}\ldots,\beta_{s}\}$ such that $\|x-\bigvee_{i=1}^{r}P_{\alpha_{i}}(x)\|<\frac{\varepsilon}{2\|y\|}$ and $\|y-\bigvee_{j=1}^{s}Q_{\beta_{j}}(y)\|<\frac{\varepsilon}{2\|x\|}$. Note that $\||\bigvee_{i=1}^{r}P_{\alpha_{i}}(x)|\|\leq\|\bigvee_{i=1}^{r}|P_{\alpha_{i}}(x)|\|=\|\bigvee_{i=1}^{r}P_{\alpha_{i}}(|x|)\|\leq\|x\|$. Therefore,
$\|x\otimes y-\bigvee_{i=1}^{r}\bigvee_{j=1}^{s}P_{\alpha_{i}}(x)\otimes Q_{\beta_{j}}(y)\|=\|x\otimes y-\bigvee_{i=1}^{r}\bigvee_{j=1}^{s}P_{\alpha_{i}}(x)\otimes Q_{\beta_{j}}+\bigvee_{i=1}^{r}P_{\alpha_{i}}(x)\otimes y-\bigvee_{i=1}^{r}P_{\alpha_{i}}(x)\otimes y\|\leq$
$\|x-\bigvee_{i=1}^{r}P_{\alpha_{i}}(x)\|\|y\|+\||\bigvee_{i=1}^{r}P_{\alpha_{i}}(x)|\|\|\bigvee_{j=1}^{s}(y-Q_{\beta_{j}}(y))\|<\varepsilon.$
So, we conclude that for each $v\in E\otimes F$, we can find indices $\{\alpha_{1},\ldots,\alpha_{n}\}$ and $\{\beta_{1},\ldots,\beta_{m}\}$ such that $\|v-\bigvee_{i=1}^{n}\bigvee_{j=1}^{m}(P_{\alpha_{i}}\otimes Q_{\beta_{j}})(v)\|<\frac{\varepsilon}{2}$.
Now, we show that for each $u\in E\widehat{\otimes}F$, we have $\|u-\bigvee_{i=1}^{n}\bigvee_{j=1}^{m}(P_{\alpha_{i}}\otimes Q_{\beta_{j}})(u)\|<\varepsilon$. This completes the proof. By density, there exists $v\in E\otimes F$ with $\|u-v\|<\frac{\varepsilon}{2}$. By the former case, $\|v-\bigvee_{i=1}^{n}\bigvee_{j=1}^{m}(P_{\alpha_{i}}\otimes Q_{\beta_{j}})(v)\|<\frac{\varepsilon}{2}$. We have
$\|u-\bigvee_{i=1}^{n}\bigvee_{j=1}^{m}(P_{\alpha_{i}}\otimes Q_{\beta_{j}})(u)\|\leq\|(u-v)-\bigvee_{i=1}^{n}\bigvee_{j=1}^{m}(P_{\alpha_{i}}\otimes Q_{\beta_{j}})(u-v)\|+\|v-\bigvee_{i=1}^{n}\bigvee_{j=1}^{m}(P_{\alpha_{i}}\otimes Q_{\beta_{j}})(v)\|<\varepsilon.$
Note that an order continuous normed lattice need not be order complete. For example the normed lattice consisting of all step-functions in $L^{2}[0,1]$ is order continuous but not order complete. So, order continuity of $E\overline{\otimes}F$ does not imply order completeness of it, in general. Nevertheless, we can have a similar version of Lemma 10 in this setting.
Lemma 11.
Suppose $E$ and $F$ are Banach lattices such that the Fremlin tensor product $E\overline{\otimes}F$ is order continuous. Moreover, assume that $\textbf{B}=(B_{\alpha})_{\alpha\in I}$ and $\textbf{C}=(C_{\beta})_{\beta\in J}$ are dense band decompositions of $E$ and $F$, respectively. Then, the collection $\textbf{A}=\{B_{\alpha}\widehat{\otimes}C_{\beta};B_{\alpha}\in\textbf{B},C_{\beta}\in\textbf{C}\}$ forms a dense band decomposition for $E\widehat{\otimes}F$.
Proof.
The proof essentially has the same idea as the proof of Lemma 10. First, note that by [1, Theorem 3.27], we conclude that $E\widehat{\otimes}F$ is also order continuous so that both $E$ and $F$ are order continuous by [17, Lemma 12]. Therefore, every band in $E$, $F$ and $E\widehat{\otimes}F$ is a projection band. For each $\alpha\in I$ and for each $\beta\in J$, assume that $B_{\alpha}=B_{x_{\alpha}}$ and $C_{\beta}=C_{y_{\beta}}$. By [3, Theorem 4.2], $B_{\alpha}\overline{\otimes}C_{\beta}$ is an order dense vector sublattice in the band in $E\overline{\otimes}F$ generated by $x_{\alpha}\otimes y_{\beta}$, denoted by ${D}_{\alpha,\beta}$, so that it is norm dense. On the other hand, every band is norm closed so that ${D}_{\alpha,\beta}$ is also a band (closed ideal) in $E\widehat{\otimes}F$ by [1, Theorem 3.8]. Therefore, A consisting of projection bands in $E\widehat{\otimes}F$. The rest of the proof is similar to the proof of Lemma 10.
Theorem 12.
Suppose $E$ and $F$ are Banach lattices such that $E\overline{\otimes}F$ is both order continuous and order complete. Moreover, assume that $S(X)$ and $S(Y)$ are the corresponding representations of $E$ and $F$, respectively as described in Theorem 8. Then, there exists a similar representation for the Fremlin projective tensor product of $E$ and $F$ in $S(X\times Y)$, as well.
Proof.
First, note that order continuity of $E\overline{\otimes}F$ implies order continuity of $E\hat{\otimes}F$ so that order continuity of both $E$ and $F$ by [17, Lemma 12]. By considering Theorem 8, there are Hausdorff spaces $X$ and $Y$, lattice isomorphisms and also $un$-homeomorphisms $T:E\to S(X)$ and $S:F\to S(Y)$. Consider the bi-injective lattice bimorphism defined via $(x,y)\to T(x)\otimes S(y)$ from $E\times F$ into $S(X\times Y)$; it induces a lattice isomorphism $T\otimes S:E\overline{\otimes}F\to S(X\times Y)$ defined via $(T\otimes S)(x\otimes y)=T(x)\otimes S(y)$. By the assumption, $E\overline{\otimes}F$ is order complete. So,by [1, Theorem 4.31], it is order dense in $E\widehat{\otimes}F$. Therefore, $T\otimes S$ can have a unique order continuous extension from $E\widehat{\otimes}F$ into $S(X,\times Y)$ by [2, Theorem 2.32]. Therefore, $E\widehat{\otimes}F$ is an order complete order dense vector sublattice of $S(X\times Y)$ so that an ideal by [2, Theorem 2.31]. So, we can have the extended $un$-topology on $S(X\times Y)$ that make it locally solid because of order continuity of $E\widehat{\otimes}F$.
Note that the extended $un$-topology has the $\sigma$-Lebesgue property by [1, Theorem 7.49]. Also, by [13, Proposition 9.1], it satisfies the pre-Lebesgue property. So, it has a Lebesgue property by [1, Theorem 3.27]. We show that $T\otimes S$ is a $un$-homeomorphism. Suppose $(u_{\alpha})\subseteq(E\widehat{\otimes}F)_{+}$ is $un$-null. There exists an increasing sequence $(\alpha_{n})$ of indices such that $u_{\alpha_{n}}\xrightarrow{un}0$ and $u_{\alpha_{n}}\xrightarrow{uo}0$ as well by [6, Corollary 3.5]. By Corollary 6, $(T\otimes S)(u_{\alpha_{n}})\xrightarrow{uo}0$ in $S(X\times Y)$. By [13, Proposition 9.2], we see that $(T\otimes S)(u_{\alpha_{n}})\xrightarrow{un}0$. Since $un$-convergence is topological, we conclude that $(T\otimes S)(u_{\alpha})\xrightarrow{un}0$.
For the converse, assume that $(T\otimes S)(x_{\gamma})\xrightarrow{un}0$. First, assume that both $E$ and $F$ have quasi-interior points so that by Proposition 9, $E\widehat{\otimes}F$ has a quasi-interior point, as well. Moreover, the corresponding topological spaces $X$ and $Y$ can be assumed to be compact. By [13, Corollary 3.2], there exists an increasing sequence $(\gamma_{n})$ of indices such that that $(T\otimes S)(x_{\gamma_{n}})\xrightarrow{un}0$. By [13, Theorem 9.5] and by passing to a further subsequence, we may assume that $(T\otimes S)(x_{\gamma_{n}})\xrightarrow{uo}0$ in $S(X\times Y)$. By using regularity of $E\widehat{\otimes}F$ in $S(X\times Y)$ and also Lemma 2, $x_{\gamma_{n}}\xrightarrow{uo}0$ in $E\widehat{\otimes}F$ so that $x_{\gamma_{n}}\xrightarrow{un}0$. Again, since $un$-convergence is topological, we see that $x_{\gamma}\xrightarrow{un}0$. For the general case, we may use the dense band decomposition for $E\widehat{\otimes}F$ as described in Lemma 11. Note that since $E$ and $F$ are order continuous, by [14, Proposition 1.a.9], they possesses dense band decompositions $\textbf{B}=(B_{\alpha})_{\alpha\in I}$ and $\textbf{C}=(C_{\beta})_{\beta\in J}$, respectively. By Lemma 11, $\textbf{A}=\{B_{\alpha}\widehat{\otimes}C_{\beta};B_{\alpha}\in\textbf{B},C_{\beta}\in\textbf{C}\}$ forms a dense band decomposition for $E\widehat{\otimes}F$. By the former case, there are compact Hausdorff spaces $(X_{\alpha})_{\alpha\in I}$ and $(Y_{\beta})_{\beta\in J}$ and also lattice isomorphisms $T_{\alpha}\otimes S_{\beta}:B_{\alpha}\widehat{\otimes}C_{\beta}\to S(X_{\alpha}\otimes Y_{\beta})$ that are $un$-homeomorphisms. Now, we can use from a similar representation as we had for Theorem 4.
Assume that $(T\otimes S)(x_{\gamma})\xrightarrow{un}0$. Write $x_{\gamma}=(y_{\gamma}^{\alpha,\beta})$ in which $y_{\gamma}^{\alpha,\beta}\in B_{\alpha}\widehat{\otimes}C_{\beta}$. Therefore, $T_{\alpha}\otimes S_{\beta}(y_{\gamma}^{\alpha,\beta})\xrightarrow{un}0$ in $S(X_{\alpha}\times Y_{\beta})$. By the former case, $(P_{\alpha}\otimes Q_{\beta})(x_{\gamma})=y_{\gamma}^{\alpha,\beta}\xrightarrow{un}0$ in $B_{\alpha}\widehat{\otimes}C_{\beta}$. Now, by using [12, Theorem 4.12], we see that $x_{\gamma}\xrightarrow{un}0$ in $E\widehat{\otimes}F$.
On behalf of all authors, the corresponding author states that there is no conflict of interest.